OpenStudy (anonymous):

On a school map, the points (8, -4), (8, -10), (2, -4), and (2, -10) represent the four corners of a sand park. What are the coordinates of the center of the sand park? (6, -9) (5, -7) (4, -2) (8, -5)

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

It's a square, so you can just use the midpoint formula for a diagonal. Use the points: (8, -4) and (2, -10) for (x1, y1) and (x2, y2)\[midpoint = \left( \frac{ x _{1} + x _{2} }{ 2 }, \frac{ y _{1} + y _{2} }{ 2 } \right)\]

OpenStudy (anonymous):

3,-3

OpenStudy (anonymous):

sorry its 5,-7 i subtracted on accident

OpenStudy (anonymous):

Good job! You got it!

OpenStudy (anonymous):

Good luck in all of your studies and thx for the recognition! @french_dame